Title1 Title2

Shalosh B. EKHAD and Doron ZEILBERGER

∑_{x,y inZ_{k}} cos^{a}(\frac{2π}{k} · x) · cos^{b}(\frac{2π}{k} · y) · cos^{c}(\frac{2π}{k} · (x + y)).

References

Shalosh B. Ekhad, c/o D. Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: ShaloshBEkhad at gmail dot com .

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: DoronZeil at gmail dot com .